# Manifolds of mappings associated with real-valued functions spaces

### v1

@@ 4 Manifolds of $\mathcal{F}$-maps on compact manifolds
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M_{1}$ and $M_{2}$ be m-dimensional compact smooth manifolds with corners and $N$ be an $n$-dimensional smooth manifold which admits a local addition. If $\Theta:M_{1}\to M_{2}$ is a smooth diffeomorphism, then the map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners. If $N$, $L$ and $K$ are smooth manifolds which admits local additions and $f:L\times K\to N$ is a smooth map and $\gamma\in\mathcal{F}(M,L)$ is fixed, then
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners and $N$ be a $n$-dimensional smooth manifold. Then the evaluation map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a $m$-dimensional compact smooth manifold with corners and $N$ be a $n$-dimensional smooth manifold with local addition. Then, for each $q\in N$, the function $\zeta_{q}:M\to N$, $p\mapsto q$ is in $\mathcal{F}(M,N)$ and the map
@@
Let $\Sigma\colon\Omega\to N$ be a local addition for $N$ and notation as in Definition 4.1 and Remark 4.3. We have $V\subseteq\Omega$ for an open 00-neighborhood $V\subseteq T_{y}N$. Then $U_{N}:=\Sigma(V)$ is an open $y$-neighborhood in $N$ and $\psi:=\Sigma|_{V}^{U_{N}}\colon V\to U_{N}$ is a $C^{\infty}$-diffeomorphism with
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N$ be an $n$-dimensional smooth manifold which admits a local addition and let $T\mathcal{F}(M,N)$ be the tangent bundle of $\mathcal{F}(M,N)$. Since the point evaluation map $\varepsilon_{p}:\mathcal{F}(M,N)\to N$ is smooth for each $p\in M$, we have
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N$ be an $n$-dimensional smooth manifold which admits a local addition and $\gamma\in\mathcal{F}(M,N)$. Then $\Theta_{N}(v)\in\Gamma_{\mathcal{F}}(\gamma)$ for each $v\in T_{\gamma}\mathcal{F}(M,N)$ and the map
@@
Working with the geometric point of view of tangent vectors, we see that $(0,\sigma)$ is equivalent to the curve $[s\mapsto s\sigma]$. Hence, for each $p\in M$ we have
@@
is in $\mathcal{F}(M,TN)$ with $\pi_{TN}\circ\Theta_{\gamma}(v)=\gamma$, making the map $\Theta_{\gamma}$ an isomorphism of topological vector spaces. ∎
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners and $N$ be an $n$-dimensional smooth manifold which admits a local addition. Since $TN$ admits local addition and the vector bundle $\pi_{TN}:TN\to N$ is smooth, the map
$\mathcal{F}(M,\pi_{TN}):\mathcal{F}(M,TN)\to\mathcal{F}(M,N),\quad\tau\mapsto\pi_{TN}\circ\tau$
@@
$\mathcal{F}(M,\pi_{TN})^{-1}(\{\gamma\})=\Gamma_{\mathcal{F}}(\gamma).$
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N$ be an $n$-dimensional smooth manifold which admits a local addition and $\pi_{TN}:TN\to N$ its tangent bundle. Then the map
$\mathcal{F}(M,\pi_{TN}):\mathcal{F}(M,TN)\to\mathcal{F}(M,N),\quad\tau\mapsto\pi_{TN}\circ\tau$
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be a $n$-dimensional smooth manifold which admits a local addition. If $f:N_{1}\to N_{2}$ is a smooth map, then the tangent map of
@@
Example 4.21
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners and $G$ be an $n$-dimensional Lie group, then we already know that the space $\mathcal{F}(M,G)$ is a Lie group (see [15]). We will give an alternative proof of this.
@@
Instead of using the set $[0,\infty)^{m}$, it is possible to generalize all results to a good collection of open subsets $\mathcal{U}$ of a locally convex, closed subset of $\mathbb{R}^{m}$, such as half-spaces, all of $\mathbb{R}^{m}$, or a disjoint union of countably many $m$-dimensional polytopes.
5 Manifolds of $\lambda$-Hölder continuous functions
@@
Remark 5.1
Since each space $\mathcal{F}_{\lambda}(U,\mathbb{R})$ is a Banach space, the axiom (PB) is verified. Indeed, let $U$ be an open subset of $\mathbb{R}^{m}$ and $V,W\in\mathcal{U}$ such that $W$ has compact closure contained in $U$ and $\Theta:U\to V$ be a $C^{\infty}$-diffeomorphism. By relative compactness of $W$, we can consider a finite open cover of convex subsets $(W_{i})_{i=1}^{k}$ for $\overline{W}$ such that $\Theta|_{W_{i}}$ is Hölder continuous and $\eta\circ\Theta|_{W_{i}}\in\mathcal{F}_{\lambda}(W_{i},\mathbb{R})$ for each $i\in\{1,...,k\}$ and $\eta\in\mathcal{F}_{\lambda}(V,\mathbb{R})$. Therefore,the map $\mathcal{F}_{\lambda}(\Theta|_{W},\mathbb{R})$ makes sense, and by Lemma 2.7, the axiom (PB) is verified.
Remark 5.2
The axiom (GL) is also verified. In fact, if $U,V\in\mathcal{U}$ with $V\subseteq U$ and $\eta\in\mathcal{F}_{\lambda}(V,\mathbb{R})$ has compact support. Let $\tilde{\eta}:U\to\mathbb{R}$ be the map defined by extending $\eta$ by 00, then $\lVert\tilde{\eta}\lVert_{\lambda}=\lVert\eta\lVert_{\lambda}$.
Therefore, the map $e_{U,V,K}^{E}:\mathcal{F}_{K}(V,E)\to\mathcal{F}(U,E)$ makes sense, and by Lemma 2.8, the axiom (GL) is verified.
Lemma 5.3
Proof.
Remark 5.4
By Lemma 2.9 and Lemma 5.3, the axiom (MU) is verified.
@@
By Lemma 5.7, the axiom (PF) is verified.
@@
Let $M$ and $N$ be finite-dimensional smooth manifolds without boundary and $0<\lambda\leq 1$. We denote the set $BC^{0,\lambda}(M,N)$ of all functions $\gamma:M\to N$ such that for each $p\in M$, there exist the charts $\phi_{M}:U_{M}\to V_{M}$ of $M$ and $\phi_{N}:U_{N}\to V_{N}$ of $N$, such that $p\in U_{M}$, $\gamma(U_{M})\subseteq U_{N}$ and $\phi_{N}\circ\gamma\circ\phi_{M}^{-1}\in\mathcal{F}_{\lambda}(V_{M},\mathbb{R}^{n})$.
@@
Let $0<\lambda\leq 1$. For each compact manifold $M$ without boundary and smooth manifold N without boundary which admits local addition, the set $BC^{0,\lambda}(M,N)$ admits a smooth manifold structure.
@@
$BC^{0,\lambda}(M,N_{1})\to BC^{0,\lambda}(M,N_{2}),\quad\gamma\mapsto f\circ\gamma$
@@
Let $M$ be a compact smooth manifold without boundary and $N$ a smooth manifold without boundary which admits a local addition. If $0<\beta\leq\lambda\leq 1$, then ${\gamma\in BC^{0,\beta}(M,N)}$ for each $\gamma\in BC^{0,\lambda}(M,N)$. Moreover, the map
$\iota:BC^{0,\lambda}(M,N)\to BC^{0,\beta}(M,N),\quad\gamma\mapsto\gamma$
@@
Let $\gamma\in BC^{0,\lambda}(M,N)$, then for each $p\in M$, there exists the charts $\phi_{M}:U_{M}\to V_{M}$ of $M$ and $\phi_{N}:U_{N}\to V_{N}$ of $N$, such that $p\in U_{M}$, $\gamma(U_{M})\subseteq U_{N}$ and $\phi_{N}\circ\gamma\circ\phi_{M}^{-1}\in\mathcal{F}_{\lambda}(V_{M},\mathbb{R}^{n})$. For each $U\in\mathcal{U}$, it is known that for $\beta\leq\lambda$ the linear operator
@@
Therefore $\gamma\in BC^{0,\beta}(M,N)$. Now, we consider the charts $(\mathcal{U}_{\gamma},\Psi_{\gamma}^{-1})$ and $(\mathcal{U}_{\iota(\gamma)},\Psi_{\iota(\gamma)}^{-1})$ in $\gamma\in BC^{0,\lambda}(M,N)$ and $\iota(\gamma)\in BC^{0,\beta}(M,N)$ respectively, then the map
@@
$\Psi_{\iota(\gamma)}^{-1}\circ\iota\circ\Psi_{\gamma}(\sigma)=(\pi_{TN},\Sigma_{N})^{-1}\circ\Big{(}\iota(\gamma),\iota(\Sigma_{N}\circ\sigma)\Big{)}$
@@
which is continuous by Proposition 3.1 and continuity of the maps $\{I_{U}\}_{U\in\mathcal{U}}$. ∎

### v2

@@ 4 Manifolds of $\mathcal{F}$-maps on compact manifolds
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M_{1}$ and $M_{2}$ be m-dimensional compact smooth manifolds with corners and $N$ be an $n$-dimensional smooth manifold which admits a local addition. If $\Theta:M_{1}\to M_{2}$ is a smooth diffeomorphism, then the map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners. If $N$, $L$ and $K$ are smooth manifolds which admits local additions and $f:L\times K\to N$ is a smooth map and $\gamma\in\mathcal{F}(M,L)$ is fixed, then
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners and $N$ be a $n$-dimensional smooth manifold. Then the evaluation map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a $m$-dimensional compact smooth manifold with corners and $N$ be a $n$-dimensional smooth manifold with local addition. Then, for each $q\in N$, the function $\zeta_{q}:M\to N$, $p\mapsto q$ is in $\mathcal{F}(M,N)$ and the map
@@
Let $\Sigma\colon\Omega\to N$ be a local addition for $N$ and notation as in Definition 4.1 and Remark 4.3. We have $V\subseteq\Omega$ for an open 0-neighborhood $V\subseteq T_{y}N$. Then $U_{N}:=\Sigma(V)$ is an open $y$-neighborhood in $N$ and $\psi:=\Sigma|_{V}^{U_{N}}\colon V\to U_{N}$ is a $C^{\infty}$-diffeomorphism with
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N$ be an $n$-dimensional smooth manifold which admits a local addition and let $T\mathcal{F}(M,N)$ be the tangent bundle of $\mathcal{F}(M,N)$. Since the point evaluation map $\varepsilon_{p}:\mathcal{F}(M,N)\to N$ is smooth for each $p\in M$, we have
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N$ be an $n$-dimensional smooth manifold which admits a local addition and $\gamma\in\mathcal{F}(M,N)$. Then $\Theta_{N}(v)\in\Gamma_{\mathcal{F}}(\gamma)$ for each $v\in T_{\gamma}\mathcal{F}(M,N)$ and the map
@@
Working with the geometric point of view of tangent vectors, we see that $(0,\sigma)$ is equivalent to the curve $s\mapsto s\sigma]$. Hence, for each $p\in M$ we have
@@
is in $\mathcal{F}(M,TN)$ with $\pi_{N}\circ\Theta_{\gamma}(v)=\gamma$, making the map $\Theta_{\gamma}$ an isomorphism of topological vector spaces. ∎
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners and $N$ be an $n$-dimensional smooth manifold which admits a local addition. Since $TN$ admits local addition and the vector bundle $\pi_{N}:TN\to N$ is smooth, the map
$\mathcal{F}(M,\pi_{N}):\mathcal{F}(M,TN)\to\mathcal{F}(M,N),\quad\tau\mapsto\pi_{N}\circ\tau$
@@
$\mathcal{F}(M,\pi_{N})^{-1}(\{\gamma\})=\Gamma_{\mathcal{F}}(\gamma).$
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N$ be an $n$-dimensional smooth manifold which admits a local addition and $\pi_{N}:TN\to N$ its tangent bundle. Then the map
$\mathcal{F}(M,\pi_{N}):\mathcal{F}(M,TN)\to\mathcal{F}(M,N),\quad\tau\mapsto\pi_{N}\circ\tau$
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be a $n$-dimensional smooth manifold which admits a local addition. If $f:N_{1}\to N_{2}$ is a smooth map, then the tangent map of
@@
Proposition 4.21
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N$ be an $n$-dimensional smooth manifold which admits a local addition $\Sigma:\Omega\to N$. Then the map
$\mathcal{F}(M,\Sigma):\mathcal{F}(M,\Omega)\to\mathcal{F}(M,N),\quad\sigma\mapsto\Sigma\circ\sigma$
Defines a local addition on $\mathcal{F}(M,N)$.
Proof.
For the open sets $\Omega\subseteq TN$ and $\Omega^{\prime}:=(\pi_{N},\Sigma)(\Omega)\subseteq N\times N$ we define the open sets
$\mathcal{F}(M,\Omega):=\{\sigma\in\mathcal{F}(M,TN):\sigma(M)\subseteq\Omega\}$
and
$\mathcal{F}(M,\Omega^{\prime}):=\{\alpha\in\mathcal{F}(M,N\times N):\alpha(M)\subseteq\Omega^{\prime}\}.$
Let $\gamma\in\mathcal{F}(M,N)$, we define $\sigma_{\gamma}:M\to TN,p\mapsto 0_{\gamma(p)}$. Then $\sigma_{\gamma}\in\Gamma_{\mathcal{F}}(\gamma)$ and the zero-section is given by
$\mathcal{F}(M,N)\to\mathcal{F}(M,TN),\quad\sigma\mapsto\sigma_{\gamma}.$
Moreover, we see that
$\mathcal{F}(M,\Sigma)(\sigma_{\gamma})(p)=\left(\Sigma\circ\sigma_{\gamma}\right)(p)=\Sigma(0_{\gamma(p)})=\gamma(p)$
hence $\mathcal{F}(M,\Sigma)(\sigma_{\gamma})=\gamma$ for each $\gamma\in\mathcal{F}(M,N)$.
Since $(\pi_{N},\Sigma):\Omega\to\Omega^{\prime}$ is a $C^{\infty}$-diffeomorphism, by Proposition 4.7, we can define the $C^{\infty}$-diffeomorphism
$\Theta:=\mathcal{F}\left(M,(\pi_{N},\Sigma)\right):\mathcal{F}(M,\Omega)\to\mathcal{F}(M,\Omega^{\prime}),\quad\sigma\mapsto(\pi_{N},\Sigma)\circ\sigma$
with inverse given by
$\Theta^{-1}:=\mathcal{F}\left(M,(\pi_{N},\Sigma)^{-1}\right):\mathcal{F}(M,\Omega^{\prime})\to\mathcal{F}(M,\Omega),\quad\alpha\mapsto(\pi_{N},\Sigma)^{-1}\circ\alpha$
Hence $\mathcal{F}(M,\Sigma)$ is a local addition on $\mathcal{F}(M,N)$. ∎
Remark 4.22
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners and $G$ be an $n$-dimensional Lie group, then we already know that the space $\mathcal{F}(M,G)$ is a Lie group (see [15]). We will give an alternative proof of this.
@@
Instead of using the set $0,\infty)^{m}$, it is possible to generalize all results to a good collection of open subsets $\mathcal{U}$ of a locally convex, closed subset of $\mathbb{R}^{m}$, such as half-spaces, all of $\mathbb{R}^{m}$, or a disjoint union of countably many $m$-dimensional polytopes.
5 Example of Manifolds of mappings
Let $m\in\mathbb{N}$ and $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$. If $\left(\mathcal{F}(U,\mathbb{R})\right)_{U\in\mathcal{U}}$ is a family of Fréchet spaces, then by Lemma 2.9, this family verify the Globalization Axiom. Moreover, if $\left(\mathcal{F}(U,\mathbb{R})\right)_{U\in\mathcal{U}}$ verifies the following axioms:
  • (PF’)
    For all $U,V\in\mathcal{U}$ such that $V$ is relatively compact in $U$ and each smooth map
    $f:U\times\mathbb{R}\to\mathbb{R}$, we have $f_{*}(\gamma):=f\circ(\text{id}_{V},\gamma|_{V})\in\mathcal{F}(V,\mathbb{R})$ for all $\gamma\in\mathcal{F}(U,\mathbb{R})$ and the map
    $f_{*}:\mathcal{F}(U,\mathbb{R})\to\mathcal{F}(V,\mathbb{R}),\quad\gamma\mapsto f\circ(\text{id}_{V},\gamma|_{V})$
    is continuous.
  • (PB’)
    Let $U$ be an open subset of $0,\infty)^{m}$ and $V,W\in\mathcal{U}$ such that $W$ has compact closure contained in $U$. Let $\Theta:U\to V$ be a smooth diffeomorphism. Then $\gamma\circ\Theta|_{W}\in\mathcal{F}(W,E)$ for all $\gamma\in\mathcal{F}(V,E)$.
  • (MU’)
    If $U\in\mathcal{U}$ and $h\in C_{c}^{\infty}(U,\mathbb{R})$, then $h\gamma\in\mathcal{F}(U,E)$ for all $\gamma\in\mathcal{F}(U,E)$.
Then it is a family of locally convex space suitable for global analysis.
Remark 5.1
In [15], Glöckner and Tárrega show that $H^{>\frac{m}{2}}(M,G)$ can be made a Lie group, where $M$ is a compact manifold of dimension $m$ (without boundary) and $G$ a finite-dimensional Lie group. This construction coincide with the construction using families of locally convex space suitable for global analysis (see Remark 4.22).
By Krikorian’s work (see [19]) we know that the set of Hölder-continuous functions has a smooth manifold structure. In this section, we intent to construct a manifold structure for this set of mappings using the spaces of sections as modeling space.
Definition 5.2
@@ 5 Example of Manifolds of mappings
Lemma 5.4
Let $U$ be an open subset of $\mathbb{R}^{m}$ and $V,W\in\mathcal{U}$ such that $W$ has compact closure contained in $U$ and $\Theta:U\to V$ be a $C^{\infty}$-diffeomorphism. Then $\gamma\circ\Theta|_{W}\in\mathcal{F}(W,E)$ for all $\gamma\in\mathcal{F}(V,E)$.
Proof.
By relative compactness of $W$, we can consider a finite open cover of convex subsets $(W_{i})_{i=1}^{k}$ for $\overline{W}$ such that $\Theta|_{W_{i}}$ is Hölder continuous and $\eta\circ\Theta|_{W_{i}}\in\mathcal{F}_{\lambda}(W_{i},\mathbb{R})$ for each $i\in\{1,...,k\}$ and $\eta\in\mathcal{F}_{\lambda}(V,\mathbb{R})$. Therefore $\gamma\circ\Theta|_{W}\in\mathcal{F}(W,E)$. ∎
Lemma 5.5
Proof.
@@
By Lemma 5.7, the axiom (PF’) is verified.
@@
Let $M$ and $N$ be finite-dimensional smooth manifolds without boundary and $0<\lambda\leq 1$. We denote the set $C^{0,\lambda}(M,N)$ of all functions $\gamma:M\to N$ such that for each $p\in M$, there exist the charts $\phi_{M}:U_{M}\to V_{M}$ of $M$ and $\phi_{N}:U_{N}\to V_{N}$ of $N$, such that $p\in U_{M}$, $\gamma(U_{M})\subseteq U_{N}$ and $\phi_{N}\circ\gamma\circ\phi_{M}^{-1}\in\mathcal{F}_{\lambda}(V_{M},\mathbb{R}^{n})$.
@@
Let $0<\lambda\leq 1$. For each compact manifold $M$ without boundary and smooth manifold N without boundary which admits local addition, the set $C^{0,\lambda}(M,N)$ admits a smooth manifold structure.
@@
$C^{0,\lambda}(M,N_{1})\to C^{0,\lambda}(M,N_{2}),\quad\gamma\mapsto f\circ\gamma$
@@
Let $M$ be a compact smooth manifold without boundary and $N$ a smooth manifold without boundary which admits a local addition. If $0<\beta\leq\lambda\leq 1$, then $\gamma\in C^{0,\beta}(M,N)$} for each $\gamma\in C^{0,\lambda}(M,N)$. Moreover, the map
$\iota:C^{0,\lambda}(M,N)\to C^{0,\beta}(M,N),\quad\gamma\mapsto\gamma$
@@
Let $\gamma\in C^{0,\lambda}(M,N)$, then for each $p\in M$, there exists the charts $\phi_{M}:U_{M}\to V_{M}$ of $M$ and $\phi_{N}:U_{N}\to V_{N}$ of $N$, such that $p\in U_{M}$, $\gamma(U_{M})\subseteq U_{N}$ and $\phi_{N}\circ\gamma\circ\phi_{M}^{-1}\in\mathcal{F}_{\lambda}(V_{M},\mathbb{R}^{n})$. For each $U\in\mathcal{U}$, it is known that for $\beta\leq\lambda$ the linear operator
@@
Therefore $\gamma\in C^{0,\beta}(M,N)$. Now, we consider the charts $(\mathcal{U}_{\gamma},\Psi_{\gamma}^{-1})$ and $(\mathcal{U}_{\iota(\gamma)},\Psi_{\iota(\gamma)}^{-1})$ in $\gamma\in C^{0,\lambda}(M,N)$ and $\iota(\gamma)\in C^{0,\beta}(M,N)$ respectively, then the map
@@
$\Psi_{\iota(\gamma)}^{-1}\circ\iota\circ\Psi_{\gamma}(\sigma)=(\pi_{N},\Sigma_{N})^{-1}\circ\Big(\iota(\gamma),\iota(\Sigma_{N}\circ\sigma)\Big)$
@@
which is continuous by Proposition 3.1 and continuity of the maps $\{I_{U}\}_{U\in\mathcal{U}}$.
Acknowledgements: The author would like to thank Helge Glöckner for his guidance during the development of this work. The author was partially supported by the FONDECYT Grant #1241719 and by ANID and DAAD (DAAD/BecasChile 2020, ID:91762237/62190017).